# TagAlign: Improving Vision-Language Alignment with Multi-Tag Classification

### v1

@@ 4 Experiments
TagAlign CC12M 53.9 33.5 33.3 53.5 32.2 41.3 +1.9 87.9 37.6 25.3 27.5 17.3 39.1 +3.0
4.1 Datasets and Evaluation Metric
We employ mean intersection-over-union (mIoU) to evaluate the performance, which is a standard metric in zero-shot semantic segmentation.
4.2 Implementation Details
In our implementation, we utilize the Vicuna as the large language model (LLM) [5] for extracting tags. The visual encoder is based on the CLIP ViT-B/16 model. The input images are resized to a resolution of $224\times 224$ pixels, and the patch size used is $16\times 16$. We follow the modifications described in [18] to adapt the CLIP image encoder for our method. For the projector, we adopt the design proposed in [41], which consists of two gated convolution blocks. However, in our implementation, we remove the upsampling interpolations. In our framework, only the lightweight projector is trainable, while all other modules are fixed. All our models are trained on a single node equipped with 8 NVIDIA A100 GPUs, with each GPU having 80GB of memory. We train the model with a batch size of $4,096$ and a constant learning rate of 1e-3 for a total of $30,000$ iterations. Notably, the learning rate in our approach is not adjusted using any heuristic tricks such as warmup or decay schedule, thereby reducing the number of hyperparameters involved in the training process. We refer to the Supplementary for more implementation details.
@@ 4.2 Implementation Details
Methods Datasets RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val
GroupViT [40] CC12M+YFCC 7.99 6.16 10.51 8.49
6.79
10.59 10.68
MaskCLIP [47] LAION-20M 11.52 11.85 12.06 11.87
12.01
12.57 12.74
TagAlign-Obj. CC12M 15.87 17.52 15.69 16.04
17.26
16.44 18.59
+ Attri. CC12M 18.75 20.31 20.64 19.24
20.88
21.23 23.69
4.3 Main Results
Zero-shot semantic segmentation. In Table 1, we summarize the zero-shot semantic segmentation performance comparison between our method and previous methods across 10 benchmark datasets, including datasets with and without a background class. To ensure a fair comparison, we report the performance of previous methods with and without the use of post-processing techniques, such as CRF [42] or PAMR [3, 39]. Our experimental results demonstrate that our method achieves significant improvements over previous state-of-the-art (SOTA) methods, using only the CC12M dataset for training. For instance, without post-processing techniques, our method outperforms the previous SOTA method TCL [3] by an average of $3.7\%$ and $3.6\%$ on datasets with and without a background class, respectively. When post-processing techniques are applied, our method maintains its superiority over previous methods by more than $1.9\%$. These results serve as compelling evidence for the effectiveness of our approach in zero-shot semantic segmentation.
Zero-shot referring segmentation. Additionally, we evaluate the performance of our method on referring segmentation datasets. Unlike the aforementioned semantic segmentation task, which primarily focuses on segmenting a set of object classes, referring segmentation entails the segmentation of entities described in unstructured natural language text, which frequently encompasses objects with distinct attributes. The experimental results, as depicted in Table 2, clearly demonstrate the substantial performance improvement achieved by our method when attributes are incorporated, compared to the case where attributes are not utilized. These results validate the effectiveness and importance of attribute-guided learning in enhancing the performance of referring segmentation.
4.4 Ablation Study
In this section, we perform a series of ablation experiments to assess the individual impact of each technique employed in our method. The evaluation is specifically tailored to discern trends in the influence of each technique on datasets with varying numbers of classes. Our method is rigorously evaluated on ImageNet-S50 (IN50) and ImageNet-300 (IN300), providing insights into the effects across datasets with different class numbers.
@@ 4.4 Ablation Study
$K$ 2.5k 5k 7.5k 10k 25k
50K
100k
IN50 27.2 27.9 28.8 29.2 28.9
29.3
29.4
IN300 15.0 16.3 17.4 17.8 18.0
18.8
19.0
Impact of the number of object tags. In this work, we establish the tag list by selecting the most frequent $K$ tags. In Table 3, we evaluate the performance for different values of $K$ by considering only the multi-tag classification loss $\mathcal{L}_{\mathrm{Tag}}$ to observe the impact of the number of tags on performance. From the table, we have the following observations. The performance on the ImageNet-S300 dataset, which contains more classes, tends to improve with larger values of $K$. This suggests that including a larger number of tags in the tag list allows the model to capture a wider range of object classes, leading to better performance on datasets with more diverse classes. Conversely, on the ImageNet-S50 dataset, which contains fewer classes, the performance tends to be better with smaller values of $K$. This indicates that a smaller tag list, focused on a subset of the most frequent tags, is sufficient to cover the object classes present in datasets with fewer classes. Considering the trade-off between complexity and performance, we choose $K=10,000$ as the default value in our work. Increasing $K$ beyond this value leads to marginal improvements.
Table 4: Performance comparison with DiHT [26] that distills a pre-trained multi-tag classification vision models to CLIP.
@@
Method Dataset IN50 IN300
TagAlign CC12M 50.4 31.5
DiHT [26] Laion-2B 10.6 9.7
Table 4: Performance comparison with DiHT [26] that distills a pre-trained multi-tag classification vision models to CLIP.
@@
[…]
Text Parser No. of Tags IN50 IN300
LLM 20000 50.4 31.5
NLTK 20000 48.7 29.5
Tag2text [12] 4585 47.6 29.0
@@
Methods IN50 IN300
Balanced Softmax Loss 50.4 31.5
Weighted Softmax Loss 49.5 30.9
Focal Loss 48.5 31.1
Asymmetric Loss 48.8 31.7
Table 6: Ablation study on the long-tailed training loss.
Impact of the long-tailed training loss. As illustrated in Fig. 5, the extracted tags are highly imbalanced. Here we explore various loss functions aimed at mitigating tag imbalance, including balanced softmax loss [29], weighted softmax loss [46], focal loss [19], and asymmetric loss [31]. The experimental results in Table 6 demonstrate that the balanced softmax loss outperforms other alternatives, making it our default choice.
Table 7: Combining with the state-of-the-art methods in different architectures. Our muti-tag supervision is synergistic with different architectures of visual encoders.
@@
Method Architecture IN50 IN300
@@
OVSegmentor + Multi-Tags GroupViT 41.0 15.7
CLIP ViT 45.9 26.2
CLIP + Multi-Tags ViT 50.4 31.5
Table 7: Combining with the state-of-the-art methods in different architectures. Our muti-tag supervision is synergistic with different architectures of visual encoders.
Figure 6: Comparison of tags extracted with text parsing methods (i.e., NLTK and LLM). The tags in red are misidentified by the parser. ‘rattan’ and ‘planter’ should not be treated separately as two tags (first row). tags like ‘week’ and ‘november’ are not visible in the image (second row). And ‘tomato’ is absent in the corresponding image (third row).
4.5 Visualization.
In Fig. 7, we visualize the similarity maps between text features and image features obtained from CLIP, TagAlign (Obj.), and TagAlign (Obj.+Attr.). We notice that the background exhibits high activation when aligning text features with image features extracted by CLIP. The similarity maps obtained from TagAlign (Obj.) indicate the alignment between text and image features is strongly affected by object-related regions, e.g., “cars” (first row) and “men” (second row). By incorporating attribute tags, TagAlign (Obj.+Attr.) can identify a more appropriate region within the given image based on the description. The visualization results further prove the capacity of TagAlign to prompt more precise alignment between image-text encoders.
Figure 7: Visualization of the similarity maps between patch features and text features generated by different methods.
@@ 5 Conclusion
In this work, we design a simple yet effective approach to assist precise alignment of the vision-language model without requiring extra annotations. Specifically, we employ a large language model to automatically parse object tags and attribute tags from the caption. These parsed tags are utilized to supervise the training procedure with multi-tag classification loss, in cooperation with the commonly used image-text contrastive loss. Extensive experiments demonstrate that our method achieves remarkable performances on 10 semantic segmentation benchmarks and referring expression segmentation benchmark.
Supplementary Material
Contents:
[…]
1
@@ Appendix B Discussion with Related Works
In both Tag2text [12] and RAM [45], a decoder module is employed to classify commonly used tags. This module incorporates multiple cross-attention layers, where the tag features serve as queries and the patch features serve as keys/values. Through this mechanism, the decoder module produces updated tag embeddings that have interacted with the patch features. Subsequently, the tag embeddings are fed to a fully connected layer to predict the probabilities of the image belonging to different tags. In contrast, our method aggregates patch features into an image feature through average pooling, and then explicitly aligns the tag features with the image feature via contrastive loss (i.e., the multi-label classification loss). Additionally, based on the principle of multiple-instance learning, this way also explicitly aligns the tag features with the patch features.
@@ C.1 Architecture Details
We employ the Vicuna-33b [5] as our large language model (LLM) to extract tags. The prompts we have devised are depicted in Figure 1. It is important to note that the tag parsing process is carried out offline and performed only once, rendering it cost-effective in practical scenarios.
@@
messages = [ {"role":"system", "content": f”””Hello, Assistant! Please analyze the caption and identify all distinct physical objects mentioned. Follow these guidelines:
(1) Analyze the caption and identify objects specifically mentioned.
(2) If an object is a proper noun or specified by a brand or model, convert it to its generic noun form.
(3) Only list tangible and visible objects that would likely be the primary focus or clear elements within an image, excluding abstract concepts or items inferred but not directly stated.
(4) Exclude any adjectives that may describe a noun. Only the primary object nouns should be included in your response.
I will provide the caption. Your response should be in the format of ’[object1, object2, …]’. If the caption, for example, reads ’An image of two grey Maine Coon cats sitting next to a blue IKEA bowl filled with daisies on a wooden Crate & Barrel table with a Sting Ray parked in the background’, you should reply with ’[cat, bowl, daisy, table, ca r]’. ”””}
@@ C.2 Training Details
Our method is trained on a single node equipped with $8$ NVIDIA A100 GPUs, with each GPU having $80$GB of memory. We train the model with a batch size of $4,096$ and a constant learning rate of 1e-3 for a total of $30,000$ iterations. We use AdamW as the optimizer. During training, only the parameters of the projector are trainable. The model takes less than $12$ hours to train on $8$ NVIDIA A100 GPUs.
@@ C.3 Testing Details
Following common practice [40], we resize each input image to have a shorter side of size $448$ pixels during inference. As described in the main paper, our approach involves computing the cosine similarity map between patch features and text features of different classes. Subsequently, the similarity map is upsampled to match the shape of the input image. We denote the upsampled similarity map as ${\mathbf{M}}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent the spatial dimensions of the image, and $C$ is the number of classes in the testing dataset.
When the testing dataset does not include a background class, we directly employ the $\operatorname{Argmax}$ operation along the class dimension of the similarity map. This operation assigns each pixel to a specific class.
In scenarios where the testing dataset includes a background class, we first apply a scaling function to the similarity map ${\mathbf{M}}$ as follows:
$\hat{{\mathbf{M}}}=\sigma(w{\mathbf{M}}-b).$ (7)
@@
Here, $\sigma$ represents the $\operatorname{Sigmoid}$ function, and $w$ and $b$ are scalar parameters. We set $w$ and $b$ to $10$ and $2.5$, respectively, following [41]. $\hat{{\mathbf{M}}}_{i,j}\in\mathbb{R}^{C}$ denotes the probabilities of each pixel belonging to the $C$ classes. Following common practice [40], we apply a thresholding strategy to $\hat{{\mathbf{M}}}$, where a pixel is assigned to the background class if the maximum value of $\hat{{\mathbf{M}}}_{i,j}$ exceeds a certain threshold.

### v3

@@ 4.2 Implementation Details
TagAlign CC12M 57.9 34.3 35.4 55.1 35.1 43.6+4.2 88.4 38.6 25.8 27.6 17.3 39.5+3.4
4.3 Main Results
Zero-shot semantic segmentation. In Table 1, we summarize the zero-shot semantic segmentation performance comparison between our method and previous methods across 10 benchmark datasets, including datasets with and without a background class. To ensure a fair comparison, we report the performance of previous methods with and without the use of post-processing techniques, such as CRF [43] or PAMR [3, 40]. Our experimental results demonstrate that our method achieves significant improvements over previous state-of-the-art (SOTA) methods, using only the CC12M dataset for training. For instance, without post-processing techniques, our method outperforms the previous SOTA method TCL [3] by an average of $6.3\%$ and $4.1\%$ on datasets with and without a background class, respectively. When post-processing techniques are applied, our method maintains its superiority over previous methods by $4.2\%$ and $3.4\%$. These results serve as compelling evidence for the effectiveness of our approach in zero-shot semantic segmentation.
[…]
@@ 4.3 Main Results
{tblr}
cells=halign=c,valign=m, cell11=halign=c, hline3=1-9, hline2=2-9, hline1,7=1.0pt, vline2,5,8=1-6, cell11=r=2, cell12=c=3, cell15=c=3, Methods & RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val
GroupViT [41] 7.99 6.16 10.51 8.49 6.79 10.59 10.68
MaskCLIP [50] 11.52 11.85 12.06 11.87 12.01 12.57 12.74
TagAlign-Obj. 15.87 17.52 15.69 16.04 17.26 16.44 18.59
+ Attri. 18.75 20.31 20.64 19.24 20.8821.23 23.69
Zero-shot referring segmentation. Additionally, we evaluate the performance of our method on referring segmentation datasets. Unlike the aforementioned semantic segmentation task, which primarily focuses on segmenting a set of object classes, referring segmentation entails the segmentation of entities described in unstructured natural language text, which frequently encompasses objects with distinct attributes. The experimental results, as depicted in Section 4.3, clearly demonstrate the substantial performance improvement achieved by our method when attributes are incorporated, compared to the case where attributes are not utilized. These results validate the effectiveness and importance of attribute-guided learning in enhancing the performance of referring segmentation.
4.4 Ablation Study
In this section, we perform a series of ablation experiments to assess the individual impact of each technique employed in our method. In addition to the commonly used PASCAL VOC (VOC) dataset, we also evaluate our method on the ImageNet-S50 (IN50) and ImageNet-300 (IN300) datasets, providing insights into how our method performs across datasets that vary in the number of classes
@@ 4.4 Ablation Study
[…]
@@
{tblr}
cells=halign=c,valign=m, column1=halign=l, hline1,5=1.0pt, hline2,3,4=1-11, vline2=1-8, $K$ 2.5k 7.5k 10k 25k 50k 100k
VOC 45.145.045.745.145.044.8
IN50 27.2 28.8 29.2 28.9 29.3 29.4
IN300 15.0 17.4 17.8 18.0 18.8 19.0
Impact of the number of object tags. In this work, we establish the tag list by selecting the most frequent $K$ tags. As shown in Table 3, we evaluate the performance for different values of $K$ by considering only the multi-tag classification loss $\mathcal{L}_{\mathrm{Tag}}$ and have the following observations. The performance on the ImageNet-S300 dataset, which contains more classes, tends to improve with larger values of $K$. This suggests that including a larger number of tags in the tag list allows the model to capture a wider range of object classes, leading to better performance on datasets with more diverse classes. Conversely, on the ImageNet-S50 dataset, which contains fewer classes, the performance tends to be better with smaller values of $K$. This indicates that a smaller tag list, focused on a subset of the most frequent tags, is sufficient to cover the object classes present in datasets with fewer classes. Considering the trade-off between complexity and performance, we choose $K=10,000$ as the default value in our work. Increasing $K$ beyond this value leads to marginal improvements.
@@
{tblr}
cells=halign=c,valign=m, column1=halign=c, hline1,2,3,4=1-5, hline1,5=1.0pt, Text Parser No. of Tags VOC IN50 IN300
LLM 20000 55.3 52.8 34.4
NLTK 20000 51.0 49.1 31.2
Tag2text [13] 4585 50.1 48.2 32.1
@@
[…]
Table 5: Performance comparison with DiHT [27] that distills a pre-trained multi-tag classification vision model to CLIP.
{tblr}
cells=halign=c,valign=m, column1=halign=c, hline1,2,3,4=1-5, hline1,4=1.0pt, Method Dataset VOC IN50 IN300
TagAlign CC12M 55.3 52.8 34.4
DiHT [27] Laion-2B 13.5 10.6 9.7
Impact of the long-tailed training loss. As illustrated in Fig. 3(b), the extracted tags are highly imbalanced. Here we explore various loss functions aimed at mitigating tag imbalance, including balanced softmax loss [30], weighted softmax loss [47], focal loss [20], and asymmetric loss [32]. The experimental results in Table 6 demonstrate that the balanced softmax loss outperforms other alternatives, making it our default choice.
[…]
@@
{tblr}
cells=halign=c,valign=m, column1=halign=c, hline1,6=1.0pt, hline2,3,4,5=1-4, Method VOC IN50 IN300
Balanced Softmax Loss 55.3 52.8 34.4
Weighted Softmax Loss 52.9 49.4 32.7
Focal Loss 52.0 48.7 32.6
Asymmetric Loss 52.7 50.0 32.3
@@
Table 7: Combining with SOTA methods in different architectures.
{tblr}
cells=halign=c,valign=m, column1=halign=c, hline1,2,4,6=1-6, hline1,6=1.0pt, Method Architecture IN50 IN300
@@
+ Multi-Tags GroupViT 41.0 15.7
CLIP ViT 47.2 27.1
+ Multi-Tags ViT 52.8 34.4
4.5 Visualization
In Fig. 5, we visualize the similarity maps between text features and image features obtained from CLIP, TagAlign (Obj.), TagAlign (Obj.+Attr.), and existing language-superivsed semantic segmentation methods, i.e. MaskCLIP [48], TCL [3], and SimSeg [43]. We notice that the background exhibits high activation when aligning text features with image features extracted by CLIP and existing methods. The similarity maps obtained from TagAlign (Obj.) indicate the alignment between text and image features is strongly affected by object-related regions, e.g., “car” (first row) and “man” (second row). By incorporating attribute tags, TagAlign (Obj.+Attr.) can identify a more appropriate region within the given image based on the description. In addition, we visualize the similarity maps of previous methods including MaskCLIP [50], TCL [3] and SimSeg [43]. The comparison reveals that the qualitative results of TagAlign (Obj.+Attr.) surpass those obtained from the aforementioned methods. The visualization results further prove the capacity of TagAlign to prompt more precise alignment between image-text encoders.
Figure 4: Comparison of tags extracted with text parsing methods (i.e., NLTK and LLM). The tags in red are misidentified by the parser. Best viewed in color.
Figure 5: Visualization of the similarity maps between patch features and text features generated by different methods. Best viewed in color.
@@ 5 Conclusion
In this work, we design a simple yet effective approach to assist precise alignment of the vision-language model without requiring extra annotations. Specifically, we employ a large language model to automatically parse object tags and attribute tags from the caption. These parsed tags are utilized to supervise the training procedure with multi-tag classification loss, in cooperation with the commonly used image-text contrastive loss. Extensive experiments demonstrate that our method achieves remarkable performances on 10 semantic segmentation benchmarks and 3 referring expression segmentation benchmarks. Besides, visual examples illustrate that integrating object and attribute supervision significantly improves the localization of attribute-specified objects. This inspires our future direction: delving into a wider variety of tag types (e.g., object relation) to further boost the vision-language alignment.
@@ Appendix B Discussion with Related Works
Specifically, in both Tag2text [13] and RAM [46], a decoder module is employed to classify commonly used tags. This module incorporates multiple cross-attention layers, where the tag features serve as queries and the patch features serve as keys/values. Through this mechanism, the decoder module produces updated tag embeddings that have interacted with the patch features. Subsequently, the tag embeddings are fed to a fully connected layer to predict the probabilities of the image belonging to different tags. In contrast, our method aggregates patch features into an image feature through average pooling, and then explicitly aligns the tag features with the image feature via contrastive loss (i.e., the multi-label classification loss). Additionally, based on the principle of multiple-instance learning, this way also explicitly aligns the tag features with the patch features.
@@ C.1 Architecture Details
We employ the Vicuna-33b [5] as our large language model (LLM) to extract tags. The prompts we have devised are depicted in Table 8. It is important to note that the tag parsing process is carried out offline and performed only once, rendering it cost-effective in practical scenarios.
@@
messages = [ {"role":"system", "content": f”””You are a Natural Language Processing (NLP) expert. I will provide you with a caption that describes a image. Please analyze the caption and identify all distinct physical objects mentioned and specific attributes. When you identify objects, please follow these guidelines: (1) Analyze the caption and identify objects specifically mentioned. (2) If an object is a proper noun or specified by a brand or model, convert it to its generic noun form. (3) Only list tangible and visible objects that would likely be the primary focus or clear elements within an image, excluding abstract concepts or items inferred but not directly stated. (4) Exclude any adjectives that may describe a noun. Only the primary object nouns should be included in your response. Based on the identified objects, please analyze the specific attribute of them following these guidelines: (5) Analyze the attribute of the identified objects specifically mentioned. (6) If one attribute consists of multiple words, shorten it to one word. I will provide the caption. Your response should be in the format of “objects: [object1, object2, …], attributes:[(attribute1, object1), (attribute2, object2),…]”. If there is no attribute in the caption, Your response should be in the format of “objects: [object1, object2, …], attributes:[]”. If the caption, for example, reads “An image of two grey Maine Coon cats sitting next to a blue IKEA bowl filled with daisies on a wooden Crate & Barrel table with a Sting Ray parked in the background”, you should reply with “objects: [cat, bowl, daisy, table, car], attributes: [(grey, cat), (two, cat), (blue, bowl), (wooden, table)]”. ”””}]
@@ C.2 Training Details
Our method is trained on a single node equipped with $8$ NVIDIA A100 GPUs, with each GPU having $80$GB of memory. We train the model with a batch size of $4,096$ and a constant learning rate of 1e-3 for a total of $30,000$ iterations. We use AdamW as the optimizer. During training, only the parameters of the projector are trainable. The model takes less than $12$ hours to train in total.
@@ C.3 Testing Details
As described in the main paper, our approach involves computing the cosine similarity map between patch features and text features of different classes. Subsequently, the similarity map is upsampled to match the shape of the input image. We denote the upsampled similarity map as ${\mathbf{M}}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent the spatial dimensions of the image, and $C$ is the number of classes in the testing dataset.
In our implementation, we calculate two similarity maps, referred to as ${\mathbf{M}}_{\text{TagAlign}}$ and ${\mathbf{M}}_{\text{CLIP}}$. The first one, ${\mathbf{M}}_{\text{TagAlign}}$, uses the patch features after the projector, i.e., the features produced by our TagAlign model. The second one, ${\mathbf{M}}_{\text{CLIP}}$, is based on the patch features before the projector, i.e., the features produced by the original CLIP model. Following [3], we merge ${\mathbf{M}}_{\text{TagAlign}}$ and ${\mathbf{M}}_{\text{CLIP}}$ using a weighted sum to obtain the final similarity map ${\mathbf{M}}$:
${\mathbf{M}}=\alpha{\mathbf{M}}_{\text{TagAlign}}+(1-\alpha){\mathbf{M}}_{\text{CLIP}}$ (7)
where $\alpha$ is the pre-defined weight.
Afterward, we apply $\operatorname{Softmax}$ over the similarity map ${\mathbf{M}}$ along the class dimension:
$\hat{{\mathbf{M}}}=\operatorname{Softmax}(s{\mathbf{M}}),$ (8)
@@
where $s$ is a scaling factor. $\hat{{\mathbf{M}}}_{i,j}\in\mathbb{R}^{C}$ denotes the probabilities of each pixel belonging to the $C$ classes. The $\operatorname{Argmax}$ opration is applied over $\hat{{\mathbf{M}}}$ for assigning each pixel to a specific class. In scenarios where the testing dataset includes a background class, we apply a thresholding strategy to $\hat{{\mathbf{M}}}$, where a pixel is assigned to the background class if the maximum value of $\hat{{\mathbf{M}}}_{i,j}$ is smaller than a certain threshold.
Following common practice [41], we resize each input image to have a shorter side of size $448$ pixels during inference. In addition, sliding windows are cropped from the image with a fixed stride and the segmentation results of different sliding windows are fused together.
As stated above, there are two inference techniques used to enhance the performance: 1) combining with CLIP; 2) using sliding windows over the image. The experiments detailed in Table 9 are designed to assess the impact of these techniques. Additionally, the previous best performance of other methods is listed in the second row for comparison. As can be seen, while these techniques do enhance our method’s performance, they are not the primary factors contributing to superiority over other methods.
[…]
Table 9: Ablation study on the inference techniques.
{tblr}
cells=halign=c,valign=m, column1=halign=l, hline1,6=1.0pt, hline2,3=1-4, vline2=1-5, Method VOC IN50 IN300
second competitor 53.8 47.4 26.6
plain TagAlign 54.5 51.1 33.0
+ combine with CLIP 54.5 51.6 33.2
+ sliding windows $\rightarrow$ TagAlign 55.3 52.8 34.4
Appendix D Visualization Examples
In addition to the visualized examples provided in the main manuscript, we present more examples in Fig. 6 to further reinforce the reliability of our observations. These additional examples demonstrate consistent findings that align with the observations made in the main manuscript.
Specifically, we find that TagAlign (Obj.) and TagAlign (Obj.+Attr.) outperform CLIP in terms of segmenting noun-based objects (e.g., “Girl”, “Line”). Furthermore, we observe that TagAlign (Obj.+Attr.) exhibits superior performance compared to TagAlign (Obj.) in segmenting objects specified by the attributes (e.g., “Smiling Girl”, “Yellow Line’‘).
These additional visualized examples provide further support for our observations, thereby reinforcing the effectiveness of TagAlign (Obj.) and TagAlign (Obj.+Attr.) in accurately aligning text features and image features compared to CLIP.
Figure 6: Visualized examples for different methods. “CLIP” denotes the original CLIP model. “TagAlign(Obj.)” indicates the utilization of object tags, while “TagAlign(Obj.+Attr.)” incorporates both object tags and attribute tags.
Figure 7: Samples of failure cases. The presented examples illustrate instances where both CLIP and TagAlign encounter challenges in accurately segmenting the given queries (captions).
Appendix E Limitation and Future Work
In this section, we present several examples that demonstrate failure cases of our method, as depicted in Fig. 7. Based on our analysis, we attribute these failures to two main reasons: 1) Inaccurate understanding of the text: In certain cases, the model exhibits difficulties in accurately comprehending the textual queries. For instance, when presented with the query “Baby Head”, the model tends to interpret both “Baby” and ”Head” as separate queries, resulting in the highlighting of both the regions corresponding to “Baby” and “Head”. Moreover, the model struggles to understand queries such as ”Person on right, not on horse”, leading to the simultaneous highlighting of both the “person” and the “horse”. 2) Inaccurate understanding of object relations: The model faces challenges in comprehending complex object relations, as exemplified by its failure to correctly understand queries like “Person in blue sweatshirt just right of the red umbrella.”
Based on the identified failure cases, we outline potential directions for future research: 1) Integration with a powerful large language model: We plan to connect our visual encoder with a large language model that possesses a better understanding of textual input. This integration can potentially enhance the model’s ability to accurately comprehend and interpret complex queries. 2) Modeling object relations: In addition to capturing objects and attributes, we intend to explore the modeling of relationships between objects. By incorporating explicit modeling of object relations, we expect to improve the ability of the model to understand and reason about complex spatial and semantic relationships among objects.